### v1

# Cross-Modal Learning for Anomaly Detection in Fused Magnesium Smelting Process: Methodology and Benchmark

Abstract
Fused Magnesium Furnace (FMF) is a crucial industrial equipment in the production of magnesia, and anomaly detection plays a pivotal role in ensuring its efficient, stable, and secure operation. Existing anomaly detection methods primarily focus on analyzing dominant anomalies using the process variables (such as arc current) or constructing neural networks based on abnormal visual features, while overlooking the intrinsic correlation of cross-modal information. This paper proposes a cross-modal Transformer (dubbed FmFormer), designed to facilitate anomaly detection in fused magnesium smelting processes by exploring the correlation between visual features (video) and process variables (current). Our approach introduces a novel tokenization paradigm to effectively bridge the substantial dimensionality gap between the 3D video modality and the 1D current modality in a multiscale manner, enabling a hierarchical reconstruction of pixel-level anomaly detection. Subsequently, the FmFormer leverages self-attention to learn internal features within each modality and bidirectional cross-attention to capture correlations across modalities. By decoding the bidirectional correlation features, we obtain the final detection result and even locate the specific anomaly region. To validate the effectiveness of the proposed method, we also present a pioneering cross-modal benchmark of the fused magnesium smelting process, featuring synchronously acquired video and current data for over 2.2 million samples. Leveraging cross-modal learning, the proposed FmFormer achieves state-of-the-art performance in detecting anomalies, particularly under extreme interferences such as current fluctuations and visual occlusion caused by heavy water mist. The presented methodology and benchmark may be applicable to other industrial applications with some amendments. The benchmark will be released at https://github.com/GaochangWu/FMF-Benchmark.
Index Terms:
Cross-modal learning, anomaly detection, Transformer, fused magnesium furnace.
I Introduction
Fused magnesia possesses numerous properties such as high temperature resistance, strong oxidation resistance and corrosion resistance. It serves as an indispensable resource in clinical surgery, aerospace, industry and other fields [1]. Fused Magnesium Furnace (FMF) is the main equipment for the production of fused magnesia, which melts powdered ore by generating a molten pool around 2850C through electric arc. Since the entire production is a continuous alternation of ore feeding and smelting, coupled with dynamic fluctuations in ore properties and non-optimal settings of smelting current [2], the temperature of molten pool can become unstable, leading to abnormal conditions in the FMF. Typical anomalies, especially the semi-molten condition [3], are detrimental to product quality and even threaten production security if not resolved in time. Therefore, timely and accurate anomaly detection is essential to the high quality, stable, and secure production of the fused magnesium smelting process.
In consideration of the intrinsic formation mechanism of anomaly, early studies focus on the detection, diagnosis or identification of the smelting current anomaly using knowledge-based or data-driven approaches. For instance, Wu et al. [3] introduced an abnormal condition identification method by constructing expert rules based on the features of current tracking error, current change rate, arc resistance, etc. Zhang et al. [4] combined multiscale kernel principal component analysis and multiscale kernel partial least analysis to extract the dominant anomaly information in a smelting process. Wang et al. [5] proposed a abnormal variable isolation method by projecting the main variables including three-phase current and voltage into a structure preserving space. However, the high-frequency fluctuation of the smelting current, stemming from variations in resistance and the unstable arcing distance induced by liquid tumbling, makes the accurate and reliable detection exceedingly challenging.
Figure 1: Cross-modal information is exploited to detect abnormal condition in fused magnesium smelting processes, as illustrated in (a). The picture at the bottom left shows an anomaly region on the furnace shell, whose visual feature is difficult to detect due to interference from heavy water mist. A novel FMF Transformer (FmFormer) is proposed using synchronous acquired video and current data, to explore the internal features of each modality by self-attention and the correlation feature across modalities by cross-attention, as shown in (b).
With the success of deep learning in artificial intelligence [6, 7], recent researches are stepping towards deep learning-based anomaly detection using image input [8, 9] or video [10, 11] acquired from industrial cameras. Comparing with smelting current, abnormal conditions of FMF demonstrate prominent visual features. Taking the semi-molten condition as an example, in its initial phase, undesired variation in the molten pool temperature is accompanied by a red dot in the local area of the furnace shell [10]. Based on this visual feature, Wu et al. [10] separated the anomaly detection task into the spatial feature extraction from 3D video using a 2D Convolutional Neural Network (CNN), and the temporal feature extraction and prediction using a recurrent neural network, i.e., Long-Short Term Memory (LSTM) [12]. Recently, Liu et al. [11] further explored a 3D convolutional LSTM [13] to learn spatio-temporal video data simultaneously. Although the vision-based anomaly detection provides a scheme with higher accuracy and better prediction consistency than the current-based detection, there are still challenges in such complex industrial environment due to visual interference (e.g., water mist and light impact from furnace flame, as shown in the bottom left image of Fig. 1(a)), which diminish detection fidelity of vision-based approaches. Therefore, it is crucial to utilize both video and current information for a comprehensive anomaly detection in fused magnesium smelting processes.
To enhance anomaly detection in fused magnesium smelting processes, we attempt, in this paper, feature exploration from both visual and current information through the cross-modal bridge. Inspired by the scaling success of Transformer models in natural language processing [14, 15] and computer vision [16, 17, 18, 19], we propose a novel cross-modal Transformer, dubbed FmFormer, to explore both spatial (visual) and temporal (visual and current) features, as illustrated in Fig. 1(b). First, we present a novel multiscale tokenization paradigm that effectively samples 3D patches with varying local receptive fields from visual modality (video) and 1D vectors from current modality, and embeds the samples into features (also known as tokens). Despite the considerably large dimensionality gap between 3D video and 1D current data, the tokenization paradigm seamlessly converts the multi-modal inputs into features with equivalent dimensions, enabling efficient exploration of intrinsically correlated features using succeeding attention mechanisms. More importantly, the proposed multiscale tokenization paradigm facilitates pixel-level anomaly detection, progressing from coarse to fine through hierarchical reassembly of the multiscale tokens. After tokenization, the proposed FmFormer processes the internal features of each modality separately through a self-attention. Subsequently, a bidirectional cross-attention is utilized to obtain correlation features in both current-to-visual direction and visual-to-current direction. By leveraging the cross-modal learning, high-fidelity anomaly detection can be achieved by decoding correlated tokens for class-level predictions or assembling current-to-visual tokens to locate anomaly regions (pixel-level predictions).
To demonstrate the effectiveness of the proposed FmFormer for learning the cross-modal information, we present a novel cross-modal benchmark for the fused magnesium smelting process. We collected over 1,000 hours of synchronously acquired videos and three-phase alternating current data from different production batches, and selected over $2.2\times 10^{6}$ samples to build the benchmark. By taking full advantage of the information from the two modalities, the proposed FmFormer is able to accurately detect the anomalies in fused magnesium smelting processes under extreme interferences, such as high frequency current fluctuation and visual occlusion caused by heavy water mist.
Summarized below are the main contributions of this paper:
  • A novel cross-modal Transformer dubbed FmFormer, for both class-level and pixel-level anomaly detection in fused magnesium smelting processes. The FmFormer progressively encodes the internal features of each modality and the correlation feature across the modalities through a cascading structure of alternatively stacked self-attention and cross-attention.
  • A novel multiscale tokenization paradigm for generating token sets with varying local receptive fields to facilitate the hierarchical reconstruction of pixel-level anomaly predictions with high localization accuracy.
  • A pioneering cross-modal benchmark with over 2.2 million samples of synchronously acquired video and current for anomaly detection in a real industrial scenario. To the best of our knowledge, the presented benchmark is the first cross-modal benchmark for anomaly detection of fused magnesium smelting processes. We are releasing the proposed benchmark at https://github.com/GaochangWu/FMF-Benchmark to foster research on cross-modal learning for industrial scenarios.
The rest of the paper is organized as follows. Section II reviews previous works related to the proposed method, including anomaly detection and cross-modal learning. Section III provides a detailed description of our proposed FmFormer. Section IV introduces the proposed cross-modal benchmark, the training recipe, and the implementation details for the FmFormer. Finally, we evaluate our method on the presented cross-modal benchmark and perform ablation studies of the proposed modules in Section V. Section VI provides conclusion and limitations of this work.
II Related Work
II-A Learning-based Anomaly Detection
Anomaly detection is defined as discovering patterns in data that do not match desired behavior, which can be mainly categorized into learning-based, statistical-based, information theory-based and spectral theory-based [20]. In this paper, we only focus on learning-based anomaly detection approaches. For time-series inputs, Yin et al. [21] converted the one-dimensional data into two-dimensional data using a sliding window scheme and implemented anomaly detection by combining convolutional layers and LSTM cells into an autoencoder architecture. Barz et al. [22] introduced a maximizes divergent intervals framework for spatio-temporal anomaly detection in an unsupervised learning manner. The proposed framework detects suspicious blocks generated by a different mechanism by modeling the probability density of data sub-blocks. To achieve unsupervised learning from only normal data, Liu et al. [23] introduced generative adversarial networks that learn to reconstruct the time-series signals in a low-dimensional space. When the reconstruction error is large, the input signal is considered to be abnormal. Due to the effective reduction of annotation cost, this unsupervised learning style has also been extended to image and video anomaly detection. However, significant noise and occlusion interference in harsh industrial environments greatly degrade model performance. For video anomaly detection with supervised learning, Zaheer et al. [24] introduced a generative adversarial learning method using 3D convolutional backbones. But convolution operation has a limited receptive field, resulting in biased local information.
To solve the local-bias problem mentioned above, the self-attention mechanism in Transformer model [14] is designed to compute the correlation of each element to all the other elements, resulting in a global perception. Based on this structure, Xu et al. [25] introduced an anomaly Transformer model that is able to uniformly model local and global information of time-series data, which is challenging for a convolutional architecture. Chen et al. [26] proposed a bidirectional spatio-temporal Transformer to predict urban traffic flow using graph-based traffic representation. Cross-attentions of past-to-present and future-to-present directions are designed to learn the temporal tendency from the traffic data. To solve high-dimensional vision tasks, Dosovitskiy et al. [16] extended Transformer by decomposing an image into a sequence of vectorized patches and converting them into token representations. This extended version is called ViT. For visual anomaly detection using 3D video input, Li et al. [27] applied a convolutional Transformer to encode 2D image slices into feature vectors, and then used another convolutional Transformer to decode these vectors into detection result.
Recently, a growing number of studies have shown that pure Transformer structures are also capable of learning representative features from high-dimensional video without using convolution-based encoders. For instance, Arnab et al. [28] extended the 2D tokenization of ViT [16] into 3D space via 3D convolution, called tubelet embedding. 3D patches are extracted from the input video to construct vectorized tokens, which are further fed into a standard Transformer. Piergiovanni et al. [29] generalized this idea and proposed to utilize tubes of different shapes to sparsely sample the input video. Since 3D patches and 2D patches are jointly extracted in the tokenization, both 3D video datasets and large-scale of 2D image datasets can be seamlessly applied for network training. Different from the aforementioned tokenization paradigms that samples 3D patches from the input video, we highlight the idea of dilated convolution [30] and propose a novel dilated tokenization to construct tokens with a larger local receptive field. This tokenization paradigm naturally forms a multiscale mechanism for efficiently reconstructing pixel-level anomaly detection.
II-B Cross-Modal Learning
Human-beings inherently have the ability to perceive and process cross-modal information such as language, sound, image, etc. With the rapid increase in the diversity of acquired information and the improvement of computing power, cross-modal learning is becoming an emerging direction in the field of artificial intelligence, e.g., hybrid imaging [31, 32, 33, 34], visual question answering [35, 36, 37], visual-text retrieval [38, 39, 40, 41], robot perception [42, 43] and chatbot (the ChatGPT [44, 15]).
A straightforward multi-modal learning scheme is to process each modality input with a different branch of network and then merge them to generate the fused feature. For instance, Zhou et al. [8] introduced a multi-source information fusion method that applies a CNN-based image recognition branch and a current processing branch. In their study, the detection results from the multi-source are individually predicted by each branch and then fused together with a support vector machine. Bu et al. [45] further applied a CNN and a Multi-Layer Perceptron (MLP) to extract the image feature and current feature, respectively, and then used these features to predict anomalies. However, the above linear fusion scheme neglects the correlation between the modalities, resulting in the underutilization of multi-modal information. In comparison, the most significant difference between the proposed FmFormer and the aforementioned studies is that we explicitly model the information interaction between video and current by designing the cross-attention to explore the bidirectional correlation between these two modalities, i.e., cross-modal learning [46].
Figure 2: An overview of the proposed FmFormer for anomaly detection in fused magnesium smelting processes. (a) The tokenization module converts 3D video and 1D current into two set of vectors, called tokens, embedded with positional information. We propose a novel multiscale video tokenization paradigm based on dilated sampling to obtain tokens with different spatial scales while keeping them with the same dimension. (b) The cross-modal Transformer encoder learns cross-modal representation from video tokens and current tokens by exploring internal features with self-attention and correlation features with bidirectional cross-attention. The bidirectional cross-attention integrates the features from the two modalities, generating current-to-visual tokens and visual-to-current tokens. (c) A multi-head decoder is developed to perform dense prediction and class prediction of abnormal conditions. The dense prediction head serves to spot anomaly regions by reassemble multi-scale video tokens (in light green) into image-like feature representations. The classification head fuses the class tokens from video and current inputs (in dark green and dark violet) into a classification vector.
Explicit cross-modal learning involves modeling the correlation between two modalities in the embedding space. For example, Ben-younes et al. [35] model the visual question answering task as a bilinear interaction between visual and linguistic features, and introduced a Tucker decomposition of the correlation tensor to explicitly control the model complexity. To reduce the dimensionality gap between 3D video and 1D text, Lei et al. [47] employed a ResNet [48] to encode 3D video into a set of 2D features and squeezed the temporal dimension of visual features via an average-pooling. Video tokens and text tokens are then concatenated and fed into a Transformer encoder. Instead of compressing temporal features, Yang et al. [49] proposed to embed each video frame into a visual token via a well-pre-trained visual-language model [50]. Wu et al. [51, 52] revisited the video classification task and converted it to a cross-modal learning problem by constructing a correlation matrix for video and language embeddings. This video classification paradigm is able to leverage well-pre-trained language models to generate precise semantic information. Although the spatio-temporal correlation is explored in these studies, each sample (e.g., video frame) in the time dimension is treated equally. Wu et al. [37] further introduced a temporal concept spotting mechanism using video and language embeddings to produce a category dependent temporal saliency map, which helps to enhance video recognition. Following these researches, we explicitly learn the correlation between different modalities through cross-attention mechanism in the embedding space. The main difference is that we employ a cascade structure with alternatively stacked self-attention and cross-attention, where the former is used for internal feature exploration for each modality.
III FmFormer for Anomaly Detection
The proposed FMF Transformer (FmFormer) has two sources of inputs, video with two spatial dimensions and one temporal dimension and three-phase alternating current with one temporal dimension. The FmFormer is composed of a tokenization module, a cross-modal Transformer encoder, and a multi-head decoder. The tokenization module converts 3D video and 1D current into two sets of 1D features. The encoder then explores the correlation between two modalities by using these features. The decoder consists of two functionally different parts, a classification head for class-level anomaly detection and a dense prediction head for spotting anomaly regions, i.e., pixel-level detection. The proposed FmFormer explores the internal features of each modality and the correlation feature across different modalities, achieving state-of-the-art performance in both anomaly detection and anomaly region spotting. The overall architecture of the proposed FmFormer is illustrated in Fig. 2.
III-A Tokenization
We prepare cross-modal features for our FmFormer by converting the high dimensional video into a set of vectorized patches and the low dimensional current into a set of vectors. Since they are one-dimensional vectors, we will call them “tokens” [14, 16, 28] throughout the rest of this paper. Fig. 2(a) illustrates the overall pipeline of the tokenization.
III-A1 Video tokenization
[…]
In addition to the standard tokenization described above, we propose a novel dilated tokenization that is capable of constructing a multiscale token set. Similar to the dilated convolution [30], we sample the element in the input video sparsely in the spatial dimensions, producing patches of the same size but with a larger local receptive field (as shown in Fig. 2(a)). For a patch $p_{v}^{d}$ of the same size of $p_{v}$, the dilated tokenization generates a smaller patch set $p_{v}^{d}\in P_{v}^{d}$ ($|P_{v}^{d}|=n_{t}\times n_{h}^{d}\times n_{w}^{d}$), where $n_{h}^{d}=\lfloor\frac{H}{(h_{v}-1)\times d+1}\rfloor$, $n_{w}^{d}=\lfloor\frac{W}{(w_{v}-1)\times d+1}\rfloor$, and $d$ is the dilation rate.
The integration of standard tokenization and dilated tokenization constructs a multiscale video patch set. Each 3D patches in $P_{v}$ and $P_{v}^{d}$ are then flattened as 1D vectors of size $1\times(t_{v}\cdot h_{v}\cdot w_{v}\cdot D_{v})$ and mapped into to tokens of the desired dimension $1\times D$. Following the vanilla Transformer [14], we prepend a learnable class token and then add positional embeddings to retain positional information. The proposed multiscale tokenization module can be formulated as:
$z_{v}=\{z_{v}^{cls},\{\phi_{v}(P^{d}_{v}),\phi_{v}(P_{v})\}\mathcal{W}_{v}^{z}\}+e_{v},$ (1)
[…]
III-A2 Current tokenization
The current sequence is denoted as $x_{c}\in\mathbb{R}^{T_{c}\times D_{c}}$, where $T_{c}$ is the sequence length of the input current and $D_{c}=3$ represents the three phases. In fact, the current sequence itself is an excellent set of tokens, we just need to map them into tokens $z_{c}$ of the desired dimension for the following Transformer layers:
$z_{c}=\{z_{c}^{cls},x_{c}\mathcal{W}_{c}^{z}\}+e_{c},$ (2)
where $\mathcal{W}^{z}_{c}\in\mathbb{R}^{D_{c}\times D}$ is the weight matrix of a linear mapping layer, $z^{cls}_{c}\in\mathbb{R}^{1\times D}$ is the concatenated current class token, and $e_{c}\in\mathbb{R}^{(T_{c}+1)\times D}$ is learnable positional embeddings. We denote the dimension of the current tokens $z_{c}\in\mathbb{R}^{N_{c}\times D}$, where $N_{c}=T_{c}+1$.
Despite the significant dimensionality gap between 3D video and 1D current, the designed tokenization paradigm is able to seamlessly convert this dimension difference to the quantity difference of tokens, while keeping any two tokens with matching dimensions. This characteristic of the tokenization contributes to elegantly migrating the attention mechanism to cross-modal learning.
III-B Cross-Modal Transformer Encoder
In the Transformer encoder, self-attention and bidirectional cross-attention are applied to encode the internal features and correlation features using video tokens $z_{v}$ and current tokens $z_{c}$. To better learn anomaly representation across modalities, we alternately cascade multiple layers of self-attention and cross-attention, as illustrated in Fig. 2(b).
III-B1 Internal feature encoding using self-attention
The tokenized representations of video and current inputs are barely aligned in the embedding space. In other words, it is impossible to find any correlation between the video tokens $z_{v}$ and the current tokens $z_{c}$. Therefore, before calculating the correlation we use self-attention mechanism to obtain the internal features of video and current tokens and roughly align them through end-to-end training.
We use Multi-Head Self-Attention (MHSA) [14] to process the video tokens $z_{v}$ and the current tokens $z_{c}$, respectively. First, the video and current tokens are mapped into queries $\mathcal{Q}$, keys $\mathcal{K}$, and values $\mathcal{V}$ with dimension $D$:
$\mathcal{Q}_{f}=\phi_{LN}(z_{f})\mathcal{W}^{\mathcal{Q}}_{f},\mathcal{K}_{f}=\phi_{LN}(z_{f})\mathcal{W}^{\mathcal{K}}_{f},\mathcal{V}_{f}=\phi_{LN}(z_{f})\mathcal{W}^{\mathcal{V}}_{f},$ (3)
with $f=v$ for video and $f=c$ for current, $\mathcal{W}^{\mathcal{Q}}_{f},\mathcal{W}^{\mathcal{K}}_{f},\mathcal{W}^{\mathcal{V}}_{f}\in\mathbb{R}^{D\times D}$ stand for the weight matrices of linear mapping layers, and $\phi_{LN}$ indicates the layer normalization [55]. Then the self-attention is formulated as:
$\phi_{A}(\mathcal{Q}_{f},\mathcal{K}_{f},\mathcal{V}_{f})=\sigma(\frac{\mathcal{Q}_{f}\mathcal{K}_{f}^{T}}{\sqrt{D_{f}}})\mathcal{V}_{f},$ (4)
where $\phi_{A}$ stands for the attention mechanism, $\sigma$ denotes the softmax non-linearity, and $D_{f}$ is the dimension of the input features. To jointly learn different representation subspaces from training instances, the MHSA concatenates multiple self-attentions as follows:
$\begin{split}\mathcal{A}_{f}&=\{(A_{f}^{(1)})^{T},\dots,(A_{f}^{(L_{s})})^{T}\}^{T}\mathcal{W}_{f}^{s},\\ A_{f}^{(l_{s})}&=\phi_{A}(\mathcal{Q}^{(l_{s})}_{f},\mathcal{K}^{(l_{s})}_{f},\mathcal{V}^{(l_{s})}_{f}),l_{s}\in\{1,\cdots,L_{s}\}\\ \mathcal{Q}_{f}&=\{(\mathcal{Q}^{(1)}_{f})^{T},\dots,(\mathcal{Q}^{(L_{s})}_{f})^{T}\}^{T},\\ \mathcal{K}_{f}&=\{(\mathcal{K}^{(1)}_{f})^{T},\dots,(\mathcal{K}^{(L_{s})}_{f})^{T}\}^{T},\\ \mathcal{V}_{f}&=\{(\mathcal{V}^{(1)}_{f})^{T},\dots,(\mathcal{V}^{(L_{s})}_{f})^{T}\}^{T},\end{split}$ (5)
where $A_{f}^{(l_{s})}$ denotes one of the self-attention head, $L_{s}$ is the number of self-attention in the MHSA, $\mathcal{W}_{f}^{s}\in\mathbb{R}^{D\times D}$ stands for a weight matrix to linearly map the concatenated self-attention features into dimension $D$, and $\mathcal{A}_{f}\in\mathbb{R}^{N_{f}\times D}$ is the resulting MHSA feature. In (5), $\mathcal{Q}_{f}$, $\mathcal{K}_{f}$ and $\mathcal{V}_{f}$ are evenly divided into $\mathcal{Q}^{(l_{s})}_{f},\mathcal{K}^{(l_{s})}_{f},\mathcal{V}^{(l_{s})}_{f}\in\mathbb{R}^{N_{f}\times\frac{D}{L_{s}}}$ for the computation of each self-attention head. The final tokens processed by the MHSA is:
$\begin{split}z^{\prime}_{f}&=\mathcal{A}_{f}+z_{f},\\ z_{f}&\leftarrow\phi_{MLP}(z^{\prime}_{f})+z^{\prime}_{f},\end{split}$ (6)
where $\phi_{MLP}$ denotes the MLP with two linear layers.
Explicitly interacting across tokens makes self-attention inherently a global operation. Therefore, the Transformer encoder is superior in the capability to capture the fine-grained features in the spatio-temporal dimensions of each modality compared to other backbones such as CNN or LSTM.
III-B2 Correlation feature encoding using bidirectional cross-attention
To facilitate feature exploration across different modalities, we utilize the cross-attention between visual and current modalities in a bidirectional manner. The implemented cross-attention has the following characteristics: i) Each token in one modality can interact with all the tokens in the other modality; ii) The resulting attention map is directional sensitive; iii) The dimension of each modality remain unchanged.
First, the video and current tokens processed by the self-attention, $z_{v}$ and $z_{c}$, are mapped into $\mathcal{Q}_{v},\mathcal{K}_{v},\mathcal{V}_{v}\in\mathbb{R}^{N_{v}\times D}$ and $\mathcal{Q}_{c},\mathcal{K}_{c},\mathcal{V}_{c}\in\mathbb{R}^{N_{c}\times D}$ using (3), respectively. Inspired by the image-language cross-modal learning in [56], we define the cross-attention in a bidirectional manner:
$\begin{split}A_{c\rightarrow v}^{(l_{c})}&=\phi_{A}(\mathcal{Q}^{(l_{c})}_{v},\mathcal{K}^{(l_{c})}_{c},\mathcal{V}^{(l_{c})}_{c})=\sigma(\frac{\mathcal{Q}^{(l_{c})}_{v}(\mathcal{K}^{(l_{c})}_{c})^{T}}{\sqrt{D/L_{c}}})\mathcal{V}^{(l_{c})}_{c},\\ A_{v\rightarrow c}^{(l_{c})}&=\phi_{A}(\mathcal{Q}^{(l_{c})}_{c},\mathcal{K}^{(l_{c})}_{v},\mathcal{V}^{(l_{c})}_{v})=\sigma(\frac{\mathcal{Q}^{(l_{c})}_{c}(\mathcal{K}^{(l_{c})}_{v})^{T}}{\sqrt{D/L_{c}}})\mathcal{V}^{(l_{c})}_{v},\\ \end{split}$ (7)
where $A^{(l_{c})}_{c\rightarrow v}\in\mathbb{R}^{N_{v}\times\frac{D}{L_{c}}}$ is the $l_{c}$th cross-attention head in current-to-visual direction, and $A^{(l_{c})}_{v\rightarrow c}\in\mathbb{R}^{N_{c}\times\frac{D}{L_{c}}}$ is the $l_{c}$th cross-attention head in visual-to-current direction. Similar to the MHSA, the Multi-Head Cross-Attention (MHCA) with the number of heads $L_{c}$ is also implemented using (5) to generate bidirectional MHCA $\mathcal{A}_{c\rightarrow v}\in\mathbb{R}^{N_{v}\times D}$ and $\mathcal{A}_{v\rightarrow c}\in\mathbb{R}^{N_{c}\times D}$. Then the resulting tokens encoded with cross-modal information are given as:
$\begin{split}z^{\prime}_{c\rightarrow v}&=\mathcal{A}_{c\rightarrow v}+z_{v},\\ z_{c\rightarrow v}&\leftarrow\phi_{MLP}(z^{\prime}_{c\rightarrow v})+z^{\prime}_{c\rightarrow v},\\ z^{\prime}_{v\rightarrow c}&=\mathcal{A}_{v\rightarrow c}+z_{c},\\ z_{v\rightarrow c}&\leftarrow\phi_{MLP}(z^{\prime}_{v\rightarrow c})+z^{\prime}_{v\rightarrow c},\end{split}$ (8)
where $z_{c\rightarrow v}$ is the current-to-visual tokens and $z_{v\rightarrow c}$ is the visual-to-current tokens.
It should be noted that the cross-attention map between the two modalities in (7) is directional sensitive, i.e., $\sigma(\frac{\mathcal{Q}^{(l_{c})}_{v}(\mathcal{K}^{(l_{c})}_{c})^{T}}{\sqrt{D/L_{c}}})\in\mathbb{R}^{N_{v}\times N_{c}}$ and $\sigma(\frac{\mathcal{Q}^{(l_{c})}_{c}(\mathcal{K}^{(l_{c})}_{v})^{T}}{\sqrt{D/L_{c}}})\in\mathbb{R}^{N_{c}\times N_{v}}$. Besides, it encodes the correlation of each token in one modality with all the other tokens in the other modality. Therefore, the bidirectional cross-attention enables the proposed FmFormer to have a global perception across modalities.
III-C Multi-Head Decoder
Since the processed tokens can be categorized into class tokens and regular tokens as described in (1) and (2), we can naturally perform multiple types of anomaly detection tasks, e.g., dense prediction (pixel-level anomaly detection) and class prediction (class-level anomaly detection), through a multi-head decoder setting, as shown in Fig. 2(c).
III-C1 Dense prediction head
The dense prediction head reassembles the multiscale video tokens into image-like feature representations [54], which are then progressively fused for spotting anomaly regions. Since the information from the current tokens and the class tokens has been integrated into the current-to-visual tokens after the Transformer encoder, we only use the regular current-to-visual tokens (i.e., ignoring the class token marked in dark green in Fig. 2(b)) for the prediction.
Let $z_{c\rightarrow v}=\{z_{c\rightarrow v}^{cls},z^{r,d}_{c\rightarrow v},z_{c\rightarrow v}^{r}\}$, where $z^{r,d}_{c\rightarrow v}\in\mathbb{R}^{(n_{t}\cdot n_{h}^{d}\cdot n_{w}^{d})\times D}$ and $z^{r}_{c\rightarrow v}\in\mathbb{R}^{(n_{t}\cdot n_{h}\cdot n_{w})\times D}$ denote the tokens stemmed from the dilated tokenization and the standard tokenization, respectively (as shown in Fig. 2(c)). First, we reassemble the multiscale video tokens by reshaping $z^{r,d}_{c\rightarrow v}$ and $z^{r}_{c\rightarrow v}$ into 3D features of sizes $n_{h}^{d}\times n_{w}^{d}\times(n_{t}\cdot D)$ and $n_{h}\times n_{w}\times(n_{t}\cdot D)$ and squeezing the channel numbers with $1\times 1$ convolutions:
$\begin{split}I_{v}^{d}&=\phi_{conv1\times 1}(\phi_{r}(z^{r,d}_{c\rightarrow v})),\\ I_{v}&=\phi_{conv1\times 1}(\phi_{r}(z^{r}_{c\rightarrow v})),\end{split}$ (9)
where $\phi_{r}$ is the reshape operation, $\phi_{conv1\times 1}$ denotes the $1\times 1$ convolution, and $I^{d}_{v}\in\mathbb{R}^{n_{h}^{d}\times n_{w}^{d}\times D}$ and $I_{v}\in\mathbb{R}^{n_{h}\times n_{w}\times D}$ are the resulting image-like features.
Note that the spatial resolutions of the two features are different, representing visual features at different local receptive field scales. We therefore blend the two features by first upsampling $I^{d}_{v}$ with a sequential stack of deconvolution (also known as transposed convolution) and convolution, and then adding them together:
$I^{b}_{v}=I_{v}+\phi_{conv}(\phi_{deconv}(I^{d}_{v})),$ (10)
where $\phi_{deconv}$ denotes the deconvolution and $I^{b}_{v}\in\mathbb{R}^{n_{h}\times n_{w}\times D}$ is the blended feature. The final dense prediction of the anomaly regions is generated by reconstructing the blended feature $I^{b}_{v}$ to the desired spatial resolution $H\times W$ with several sequential stacks of transposed convolution and convolution:
$\hat{y}^{pix}=\phi_{conv_{1\times 1}}(\phi_{conv}(\phi_{deconv}(\cdots\phi_{conv}(\phi_{deconv}(I^{b}_{v}))\cdots))),$ (11)
where $\hat{y}^{pix}\in\mathbb{R}^{H\times W\times K}$ denotes the pixel-level prediction result and $K=2$ indicates the possibility of normal and abnormal.
III-C2 Classification head
The classification head transforms the class tokens into an anomaly prediction, i.e., a classification vector. First, we map the class tokens (marked in dark green and dark violet in Fig. 2(c)) processed by the Transformer encoder, $z_{c\rightarrow v}^{cls}\in\mathbb{R}^{1\times D}$ and $z_{v\rightarrow c}^{cls}\in\mathbb{R}^{1\times D}$, to vectors $\hat{y}_{c\rightarrow v}^{cls}\in\mathbb{R}^{1\times K}$ and $\hat{y}_{v\rightarrow c}^{cls}\in\mathbb{R}^{1\times K}$ using two small MLPs. Then the detection result is obtained via a simple linear fusion mechanism formulated as follows:
$\hat{y}^{cls}=\sigma(\hat{y}_{c\rightarrow v}^{cls}+\hat{y}_{v\rightarrow c}^{cls}),$ (12)
where $\sigma$ denotes the softmax non-linearity and $\hat{y}^{cls}\in\mathbb{R}^{1\times K}$ indicates the anomaly detection result.
IV Benchmark and Network Training
IV-A Fused Magnesium Smelting Process Benchmark
The benchmark contains cross-modal data from fused magnesium smelting processes for a total of 3 production batches, in which the video data is captured by industrial cameras, and the current data is sampled by PLC control systems, as shown on the right of Fig. 1(a). Signal generators are employed to trigger industrial cameras and PLC control systems to capture video frames and currents, thus ensuring synchronous acquisition. The raw data contains a total of more than 1,000 hours of synchronized videos and current sequences. However, directly using the raw dataset as the benchmark causes class imbalance, as abnormal conditions are infrequent. Therefore, we carefully selected over $2.2\times 10^{6}$ samples (about 25 hours, 25 samples per second, and spatial resolution of $1440\times 2560$ for each raw video frame) through the entire raw dataset to keep the ratio of normal to abnormal samples close to $1:1$. The benchmark comprises two types of datasets: a pixel-level annotated dataset with approximately $2.7\times 10^{5}$ samples and a class-level annotated dataset with around $2.0\times 10^{6}$ samples. More detailed statistics for the benchmark are listed in Table I. The fused magnesium smelting process benchmark will be available at https://github.com/GaochangWu/FMF-Benchmark.
TABLE I: Statistics of the fused magnesium smelting process benchmark.
Dataset Pixel-level Class-level Total
annotated dataset annotated dataset
Normal 129,859 940,113 1,069,972
Abnormal 140,062 1,026,663 1,166,725
Total 269,921 1,966,776 2,236,697
Figure 3: Efficient data annotation based on the piecewise temporal consistency of abnormal conditions in fused magnesium smelting processes. (a) The initial annotation is achieved by the following steps: i) determining the onset, apex and offset times of anomalies according to the features summarized by domain experts; ii) labelling the sparse set of frames (keyframes) with boxes; iii) propagating boxes to other frames using their adjacent keyframes. (b) The annotation refinement is implemented by applying a weighted median filter to each label, providing a more accurate representation of the anomaly region.
The annotation of the large volume cross-modal data requires a mass of human labour. To accelerate the data annotation, we take advantage of the piecewise temporal consistency of abnormal conditions in fused magnesium smelting processes. More specifically, despite the data can be disturbed by noise or visual occlusion, the actual anomaly remains constant for a certain period of time. Based on this observation, we first determine the onset (starting), apex (highest intensity) and offset (ending) times of anomalies according to the visual and current features summarized by domain experts, and coarsely annotate a sparse set of frames using boxes, each of which is called a keyframe. A more detailed description of the normal/abnormal current features can be referred in [3, 2, 1]. The coarse labels are then propagated to other frames between each pair of adjacent keyframes, providing every frame a coarse label, as shown in Fig. 3(a). To obtain more accurate labels for the pixel-level anomaly detection, we adopt a weighted median filter [57] to refine each label guided by the corresponding frame, as shown in Fig. 3(b). With the help of the efficient data annotation, the presented benchmark provides approximately $2.7\times 10^{5}$ cross-modal samples with pixel-level labels. Fig. 4 demonstrates some of the examples of video frames (superimposed with labels) and current curves in the presented cross-modal benchmark.
Figure 4: Examples of the fused magnesium smelting process benchmark. Each example demonstrates two frames (superimposed with the corresponding pixel-level labels) of a video clip and three-phase alternating current curve. In our cross-modal benchmark, we provide about 270,000 pixel-level annotated samples and 2.1 million class-level annotated samples.
IV-B Training Recipe
Our training objective involves the minimization of a dense prediction loss and a classification loss:
$\arg\min_{\mathcal{W},e}\sum_{\langle x_{v},x_{c},y\rangle}\mathcal{L}^{pix}(\hat{y}^{pix},y)+\alpha\mathcal{L}^{cls}(\hat{y}^{cls},y),$
where $\mathcal{W}$ and $e$ are the learnable weights and learnable positional embeddings in the proposed FmFormer, $\langle x_{v},x_{c},y\rangle$ is the training set of input video $x_{v}$, input current $x_{c}$, and label $y$ triples, $\mathcal{L}^{pix}$ is the dense prediction loss term, $\mathcal{L}^{cls}$ is the classification loss term, and $\alpha$ is a hyperparameter to control the balance of the two terms. Specifically, we use the pixel-level cross-entropy loss for the dense prediction:
$\mathcal{L}^{pix}(\hat{y}^{pix},y)=-\frac{1}{HW}\sum_{h,w}\sum_{k\in K}y_{(h,w,k)}\log\frac{\exp\hat{y}^{pix}_{(h,w,k)}}{\sum_{k^{\prime}}\exp\hat{y}^{pix}_{(h,w,k^{\prime})}},$
where $(h,w,k)$ indicates an element index in the prediction result $\hat{y}^{pix}$ or the pixel-label $y$. For the classification loss, we first aggregate the pixel-level labels $y$ to the class-level label by determining whether there is anomaly, i.e., the type is classified as abnormal if the label $y$ contains any anomaly region; otherwise, it is categorized as normal. Then we use class-wise cross-entropy loss for the classification:
$\begin{split}\mathcal{L}^{cls}(\hat{y}^{cls},y)&=-\sum_{k\in K}\mathcal{G}(y)_{(k)}\log\hat{y}^{cls}_{(k)},\\ \mathcal{G}(y)&=\begin{cases}[0,1],&\text{if }\sum_{h,w}y_{(h,w,2)}>\tau\\ [1,0],&\text{otherwise},\end{cases}\end{split}$ (13)
where $\mathcal{G}(\cdot)$ denotes the aggregation operation that converts pixel-level labels into class-level labels as described above, and $\tau$ is a threshold which we empirically set to $0.5$.
IV-C Implementation Details
IV-C1 Architecture details
We use a kernel of size $2\times 8\times 8$ (time $t_{v}$, height $h_{v}$, and width $h_{w}$) and a kernel of dilation rate of 2 with the same kernel size for the video tokenization. Throughout the proposed FmFormer, each token has a dimension of $D=96$. In the cross-modal Transformer encoder, the numbers of heads in the MHSA and MHCA are set to $L_{s}=3$ and $L_{c}=3$, respectively. We alternately cascade $L_{e}=6$ layers of MHSA and MHCA as shown in Fig. 2(b). In the dense prediction head, the $1\times 1$ convolution layers in (9) have channel number $D=96$, and each transposed convolution layers in (10) and (11) has a $2\times 2$ kernel size with stride $2\times 2$. For simplicity, each convolution layer in (10) and (11) is practically a sequential connection of a $3\times 3$ convolution, a batch normalization, and a ReLU non-linearity. Due to the spatial reconstruction, we set the channel numbers of the transposed convolution layers in (11) as $[128,64,32,\cdots]$. The MLPs in the MHSA, MHCA (in (6) and (8)) and the classification head are sequential connections of a layer normalization, a linear layer, a GELU non-linearity, and another linear layer.
IV-C2 Training details
For the joint training of the dense prediction head and the classification head, we use the pixel-level annotated dataset as described in Section IV-A. Note that pixel-level labels can also be converted into class-level labels. The dataset is divided into a training set with about $2.15\times 10^{5}$ examples and a test dataset with $0.55\times 10^{5}$ examples. In both the training and test datasets, the ratio of normal examples to abnormal examples remained close to $1:1$. We crop out areas of interest (i.e., furnaces) from the raw videos for the training and testing. The proposed FmFormer is implemented by using the Pytorch framework [58]. The AdamW solver [59] is applied as the optimization method, in which the batch size is set to 64. The step learning rate decay scheme is adopted with an initial learning rate of $5\times 10^{-4}$, which then decays to $5\times 10^{-5}$ after 20 epochs. The network converges after 30 epochs of training, which takes about 5 hours on an NVIDIA TESLA V100.
V Experiments
In this section, we evaluate the proposed FmFormer on the fused magnesium smelting process benchmark and compare it with several state-of-the-art learning-based anomaly detection methods that apply both unimodal inputs (current or visual) and cross-modal inputs. In the experiments, quantitative evaluations and visual comparisons specifically for dense prediction are performed. In addition, we empirically investigate the modules in the proposed FmFormer through several ablation studies.
V-A Evaluation Metrics
We employ the following performance metrics that are commonly used in anomaly detection:
  • Accuracy ($Acc$) indicates the ratio of correctly predicted samples with respect to all the testing samples, i.e, $Acc=(TP+TN)/(TP+TN+FP+FN)$, where $TP$, $TN$, $FP$ and $FN$ means true positive, true negative, false positive and false negative, respectively.
  • F1-score ($F1$) is a comprehensive metric that considers precision $P=TP/(TP+FP)$ and recall, $R=TP/(TP+FN)$. Then the F1-score is computed as $F1=2P\cdot R/(P+R)$.
  • False Detection Rate (FDR), also known as false positive rate, is defined as the ratio of incorrectly predicted normal samples with respect to all the real normal (negative) samples, i.e., $FDR=FP/(TN+FP)$.
  • Miss Detection Rate (MDR) is also known as false negative rate that indicates the ratio of incorrectly predicted abnormal samples to all the real abnormal (positive) samples, i.e., $MDR=FN/(TP+FN)$.
  • Mean Intersection over Union (mIoU) is employed to evaluate the performance of dense prediction, which is typically used for image segmentation. It is defined as the ratio of the intersection to the union of the pixel-level ground truth and the dense prediction result, i.e., $mIoU=\frac{1}{K}\sum_{k\in K}TP/(TP+FN+FP)$.
TABLE II: Quantitative comparison with state-of-the-art methods on the fused magnesium smelting process benchmark.
Method Modality Classification Dense prediction
Visual Current $Acc\uparrow$ $F1\uparrow$ $FDR\downarrow$ $MDR\downarrow$ $mIoU(\%)\uparrow$
Expert system [3] 0.6851 0.7091 0.3586 0.1164 -
Informer [60] 0.7600 0.7960 0.3793 0.1164 -
Flowformer [61] 0.7776 0.7950 0.2625 0.1866 -
Flashformer [62] 0.7949 0.8065 0.2181 0.1935 -
iTransformer [63] 0.7164 0.7305 0.2934 0.2748 -
FmFormer (Ours) 0.7754 0.7688 0.1448 0.2954 -
3DCRNN [11] 0.9690 0.9708 0.0380 0.0247 0.8289
ViT [16] 0.9650 0.9671 0.0408 0.0298 0.8124
ViViT [28] 0.9687 0.9708 0.0466 0.0174 0.8321
TubeViT [29] 0.9715 0.9732 0.0344 0.0231 -
FmFormer (Ours) 0.9731 0.9744 0.0224 0.0307 0.8382
DCNN-SVM [8] 0.8785 0.9086 0.2259 0.0282 -
SSFGGAN [45] 0.8595 0.8967 0.2706 0.0245 -
Unicoder-VL [64] 0.9680 0.9700 0.0428 0.0224 0.8185
ClipBERT [47] 0.9757 0.9774 0.0340 0.0157 0.8174
CAPTURE [65] 0.9744 0.9758 0.0260 0.0250 0.8312
BiLM [49] 0.9763 0.9780 0.0182 0.0274 0.8360
FmFormer (Ours) 0.9836 0.9846 0.0216 0.0116 0.8409
V-B Comparison With State-of-the-Art Methods
We demonstrate the effectiveness of the proposed FmFormer by comparing it with several baseline methods using unimodal settings (current/visual) and a cross-modal setting.
V-B1 Current modality
Four state-of-the-art Transformer-based methods, including Informer [60], Flowformer [61], Flashformer [62] and iTransformer [63], are evaluated for current-based anomaly detection. Informer [60] employs a ProbSparse self-attention mechanism that generates sparse query-key pairs for efficient time-series modeling. Flowformer [61] linearizes Transformer free from specific inductive biases by applying the property of flow conservation into attention. Flashformer [62] uses tiling operation to reduce the number of GPU memory reads/writes to achieve IO-aware attention mechanism. iTransformer [63] embeds each series independently to the variate token instead of embedding temporal token in the vanilla Transformer. In addition, a classical data-driven expert system method by Wu et al. [3] is also compared.
Table II presents the quantitative comparison with the baseline methods on the proposed fused magnesium smelting process benchmark. For current modal input, the proposed FmFormer achieves comparable performance with state-of-the-art Transformer-based methods that specifically designed for time-series prediction. Nevertheless, from the performance perspective, anomaly detection using pure current information is still far from practical industrial applications.
V-B2 Visual modality
Four state-of-the-art methods or backbones are evaluated, which are 3DCRNN [11], ViT [16], ViVit [28] and TubeVit [29]. 3DCRNN [11] is a typical convolution recurrent-based anomaly detection framework specifically for fused magnesium smelting processes, utilizing 2D convolutional LSTM cells [13] to extract 3D (2D spatial and 1D temporal) features. ViT [16], ViViT [28] and TubeViT [29] are three Transformer-based methods specifically designed for video input, and employ non-overlapping 2D patches [16], 3D tubes [28] and 3D tubes of different shapes [29] for tokenization, respectively.
As shown in Table II, the visual models exploit visual features that are more stable and prominent than current information, which provide them a distinct advantage. Besides, 3DCRNN [11] achieves comparable performance to the Transformer-based methods, ViT [16], ViVit [28] and TubeVit [29], due to the effective spatial-temporal modeling of convolutional LSTM units. Among the unimodality-based methods using visual input, our FmFormer achieves the best comprehensive performance.
Figure 5: Visual comparison of the proposed FmFormer (cross-modality) with three state-of-the-art Transformer-based methods for pixel-level anomaly detection on three challenging cases. In each case, two close frames in a video are displayed to demonstrate the abnormal dynamics. We superimpose the pixel-level detection result of each method with the corresponding frame for better viewing. In the first case, intense flame light of the furnace (first frame) interferes with most detection methods. The anomaly is not detected until the light interference subsided (second frame). In the second case, visual occlusion from heavy water mist affects the compared methods, resulting in miss detection of the subtle anomaly (second frame). In the third case, fluctuations in ambient light influence the accuracy of anomaly location of the compared methods. In these challenging cases, the proposed FmFormer considers current information as a prompt for normal or abnormal conditions, thus demonstrating better robustness under disturbances and temporal consistency under occlusions.
TABLE III: Performance of the FmFormer without and with the dilated tokenization mechanism.
Modality Tokenization Classification Dense prediction
$Acc\uparrow$ $F1\uparrow$ $FDR\downarrow$ $MDR\downarrow$ $mIoU(\%)\uparrow$
Visual w/o dilated tokenization 0.9694 0.9709 0.0217 0.0382 0.8369
w dilated tokenization 0.9731 0.9744 0.0224 0.0307 0.8382
Cross-modal w/o dilated tokenization 0.9757 0.9772 0.0321 0.0171 0.8376
w dilated tokenization 0.9836 0.9846 0.0216 0.0116 0.8409
V-B3 Cross-modality
Six state-of-the-art methods are evaluated, including DCNN-SVM [8], SSFGGAN [45], Unicoder-VL [64], ClipBERT [47], CAPTURE [65] and BiLM [49]. DCNN-SVM [8] and SSFGGAN [45] achieve multi-modal learning through linear fusion of features from different modalities, which are designed specifically for anomaly detection in fused magnesium smelting processes. Unicoder-VL [64] encodes vectorized 2D image patches conjointly with 1D sequence (current) into Transformer backbones. ClipBERT [47] and BiLM [49] first employ pre-trained backbones to extract visual features from the video modality, and then embed the visual features together with 1D sequence into Transformers. The different is that ClipBERT [47] squeezes the temporal dimension of visual features via an average-pooling for efficiency. CAPTURE [65] also adopts a pipeline of self-attention and cross-attention for cross-modal learning. The main difference between our FmFormer and CAPTURE [65] is that we use a cascading structure to alternatively stack self-attention and cross-attention. This structure helps the network to capture internal features in each modality and correlation features across modalities in a progressive manner. Although most of the baseline methods are designed for classification task, we extend the Transformer-based models to dense prediction task by assembling tokens into image-like features as introduced in Section III-C1.
As indicated by the quantitative results in Table II, the Transformer-based models using cross-modal input (Unicoder-VL [64], ClipBERT [47] and BiLM [49]) generally outperform models using single visual modality (ViT [16], ViVit [28] and TubeVit [29]). Compared with conventional CNN-based methods using linear fusion (DCNN-SVM [8] and SSFGGAN [45]), Transformer-based methods (Unicoder-VL [64], ClipBERT [47] and BiLM [49]) generally demonstrates higher performance, as shown in Table II since the explicit correlation modeling achieved via Transformer encoder. Among methods using cross-modal input, the proposed FmFormer achieves superior performances, confirming its effectiveness for both class-level prediction and pixel-level prediction.
V-B4 Dense prediction
Fig. 5 visualizes the qualitative comparison of dense predictions of the proposed FmFormer with three state-of-the-art Transformer-based methods, ViViT [28] (visual), ClipBERT [47] (cross-modal), and BiLM [49] (cross-modal), on three representative cases. In the first case, intense flame light of the furnace has stronger visual saliency than the anomaly region (please see the first frame). The compared methods fail to detect the anomaly under the strong disturbance until the disturbance subsides (second frame). In the second case, heavy water mist occludes the visual features of the anomaly, so the detection method can only rely on the information from the previous frames and three-phase alternating current. In this case, ViViT [28] with only visual input fails to detect the anomaly. In the first frame, ClipBERT [47] and BiLM [49] with cross-modal input detect the anomaly successfully, but predict the wrong location. In the second frame, they also fail to spot the abnormal condition. In the third case, fluctuations of ambient light lead to variations of visual features, which in turn cause the compared methods to misestimate the anomaly regions (see ClipBERT [47] and BiLM [49]) or even produce missed detection (see ViViT [28]). In comparison, the proposed FmFormer leverages current information as a prompt for normal or abnormal conditions, thereby demonstrating robustness under disturbances and temporal consistency under occlusions in complex industrial environment.
V-C Ablation Study
In this section, we empirically analyse the proposed FmFormer by performing the following ablation studies.
Figure 6: Performance of the FmFormer under different settings of input sequence lengths (horizontal coordinate).
V-C1 Sequence length
Benefiting from the designed structure, our FmFormer can be adapted to input of variable sequence length without network retraining or fine-tuning [^1] Since we employ learnable positional embeddings for visual input and current input, bilinear interpolation is used to accommodate inputs of different sequence lengths.. It should be noted that since the input sequences of video and current represent the operating state at the current instant and how that state may change, we only need to ensure the final frame and three-phase current values in the input sequences belong to the same instant, while the sequence lengths of the two modalities do not need to be exactly the same. The length of the input sequence of each modality influences the network to perceive the dynamic features of normal or abnormal conditions.
In this experiment, we investigate the performance of our FmFormer under different settings of input sequence lengths, as shown in Fig 6. For the visual modality, the performance of the proposed method continues to increase when the video sequence length is longer than 1.5 seconds. However, the MDR tends to increase when the input sequence is too long. So we chose an input sequence length of 1.5 seconds. For the current modality, the performance of the proposed method improves as the sequence length increases since it is less influenced by the current noise. However, excessively long sequences of current input may drive the self-attention mechanism to tend to focus on the working conditions of other time series instead of the current moment, resulting in the performance degradation. In this experiment, the best performance is achieved when the length of the input current sequence is around 220 seconds.
V-C2 Multiscale tokenization
In this experiment, the effectiveness of the proposed multiscale tokenization module is studied by comparing the results without and with the dilated tokenization mechanism. Note that the multiscale feature blending formulated in (10) is also removed when the dilated tokenization mechanism is not utilized. As shown in Table III, the models using the proposed multiscale tokenization module achieve overall better performance in both single (visual) modality and cross-modal input settings, which demonstrate the effectiveness of the proposed module.
TABLE IV: Performance of the FmFormer using unimodal and cross-modal input, where “LF” indicates the simple linear fusion mechanism in (12).
Modality $Acc\uparrow$ $F1\uparrow$ $FDR\downarrow$ $MDR\downarrow$
Visual 0.9731 0.9744 0.0224 0.0308
Current 0.7754 0.7688 0.1448 0.2954
Cross-modality+LF 0.9748 0.9764 0.0354 0.0159
V-C3 Unimodality vs. cross-modality
To investigate the influence of each modality on the performance of anomaly detection, we degrade the proposed FmFormer to a unimodal method by using only one modality as input. Due to the absence of cross-modal input, we replace the MHCA with the MHSA in the Transformer encoder and remove the linear fusion mechanism in (12). Table IV lists the performance of the proposed FmFormer using a single visual or current modality. As expected, the visual model outperforms the current model because the visual features are more stable than the current features.
TABLE V: Performance of the FmFormer under different settings of modality interaction mechanisms. The first row (ID 0) serves as a baseline using the single visual modality. Since the linear fusion mechanism does not interact with the dense prediction head, the dense prediction results with/without it are the same.
ID Input Mechanism Classification Dense prediction
MHSA MHCA Linear fusion $Acc\uparrow$ $F1\uparrow$ $FDR\downarrow$ $MDR\downarrow$ $mIoU(\%)\uparrow$
0 Visual 0.9731 0.9744 0.0224 0.0307 0.8382
1 Cross-modality 0.9748 0.9764 0.0354 0.0159 0.8382
2 Cross-modality 0.9711 0.9726 0.0253 0.0318 0.7627
3 Cross-modality 0.9757 0.9769 0.0187 0.0291 0.8409
4 Cross-modality 0.9836 0.9846 0.0216 0.0116 0.8409
In this study, the performance of adopting cross-modal input is also verified by simply integrating the two models with unimodal inputs through a linear fusion mechanism in (12). Note that no further training or fine-tuning is performed for the linear fusion. As shown in Table IV, despite the poor performance of the current model, it can be integrated with the visual model to improve the comprehensive performance, especially the MDR decreased by nearly half compared to the vision model. This ablation study fully verifies that cross-modal input can effectively enhance anomaly detection.
V-C4 Effectiveness of modality interaction mechanisms
In this experiment, we validate the effectiveness of the applied modality interaction mechanisms, including the MHSA for internal features encoding within each modality, the MHCA for correlation feature exploration across modalities, and the simple linear fusion mechanism for explicitly weighting the predictions from two modalities. For the settings with only MHSA or MHCA, we replicate the specific interaction mechanism to keep the numbers of the model parameters the same.
Table V lists the performance of the proposed FmFormer using different settings of modality interaction mechanisms, where the first row (ID 0) serves as a baseline using the single visual modality. It should be noted that the linear fusion mechanism does not interact with the dense prediction head, and thus, the dense prediction results with/without it are the same (e.g., ID 0 vs. ID 1 and ID 3 vs. ID 4). As mentioned in Section V-C3, utilizing a simple linear fusion of the two modalities (ID 1) is also able to improve the model performance compared to the single modality (ID 0). While in comparison, the model with only the MHCA (ID 2) has an unsatisfactory performance and is even worse than the baseline setting using a single modal input (ID 0), especially for dense prediction tasks. It indicates that directly modeling the cross-modal correlation by using cross-attention without self-attention feature encoding degrades the model performance. The self-attention provides sufficient preparation for correlation exploration in the cross-attention. In addition, the model with the MHCA (ID 3) outperforms the model with the linear fusion mechanism (ID 1), which verifies the effectiveness of this explicit modality interaction mechanism. Among these settings, the FmFormer with the full modality interaction mechanisms (ID 4) achieves the best comprehensive performance.
VI Conclusion and Limitations
In this paper, we introduce a novel Transformer, dubbed FmFormer, through the lens of cross-modal learning to enhance anomaly detection in a scenario of fused magnesium smelting process. In the proposed FmFormer, a multiscale tokenization module is developed to handle the problem of large dimensionality gap between the two modalities of video and three-phase alternating current. On this basis, a cross-modal Transformer encoder is employed to alternatively explore the internal features of each modality and the correlation features across modalities. Through a multi-head decoder, our FmFormer is able to preform class-level anomaly prediction and pixel-level anomaly region detection. Interestingly, despite the poor detection accuracy when using a single current modality, the comprehensive performance of the model can still be improved through a simple linear fusion. Furthermore, by taking advantage of the cross-modal learning, the proposed method achieves an accurate anomaly detection under extreme interferences such as current fluctuation and visual occlusion from heavy water mist. To demonstrate the effectiveness of the FmFormer, we present the first cross-modal benchmark for anomaly detection of fused magnesium smelting processes, which possesses synchronously acquired video-current data and pixel-level labels. We hope it will be helpful to the research community of cross-modal learning in industrial scenarios.
Limitations. Promising directions not discussed in this work include joint training using a mixture of image-current cross-modality, video-current cross-modality, and unimodal inputs, which may involve adaptive sampling (tokenization) depending on the input modes. Adaptation to different input modes may help to solve the problem of temporary failure of a certain modality in practical applications. Another limitation is that only current-to-video tokens stemmed from the dilated tokenization and standard tokenization are used for the dense prediction. Despite the explicit interaction of the two modalities in the Transformer encoder, the performance of dense prediction could still be improved if video-to-current tokens and class tokens are used.